SPECTROSCOPIC APPLICATIONS OF STATE-SELECTED SYMPATHETICALLY-COOLED MOLECU-LAR IONS

\underline{XIN TONG}, MATTHIAS GERMANN and STEFAN WILLITSCH, Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

Cold molecular ions prepared by sympathetic cooling with laser-cooled ions in an ion trap^{a} represent attractive systems for new spectroscopic experiments. The long trapping times (up to hours) and state lifetimes (up to minutes)^{b} in an almost perturbation-free environment enable the long interaction times required for the study of forbidden spectroscopic transitions which have not been accessible before in molecular ions. Here, we present a proof-of-principle experiment for the investigation of dipole-forbidden infrared transitions in cold N{+ \atop 2} ions using quantum-cascade-laser technology in combination with action spectroscopy. Because sympathetic-cooling experiments typically use small ensembles of tens to hundreds of ions, the confinement of their population into a single quantum state is essential to improve the sensitivity of our experiments. This is achieved by state-selective generation of the ions using threshold photoionization followed by sympathetic cooling^{c}. Finally, we discuss the experimental requirements for performing highly sensitive spectroscopic measurements on trapped, cold molecular ions and present an outlook on current developments which employ quantum-logic methods for nondestructive spectroscopic studies on single sympathetically cold molecular ions.^{d}

aS. Willitsch, M. Bell, A. Gingell and T. P. Softley Phys. Chem. Chem. Phys. \underline{10}, 7200 2008.

bX. Tong, A. Winney and S. Willitsch Phys. Rev. Lett. \underline{105}, 143002 2010.

cX. Tong, D. Wild and S. Willitsch Phys. Rev. A \underline{83}, 023415 2011.

dJ. Mur-Petit, J. J. Carcia-Ripoll and S. Willitsch et al. Phys. Rev. A \underline{85}, 022308 2012.